# A simple, flexible method for timing cross-calibration of space missions

### v1

@@ A.8 NICER
  • DE200 ephemeris, (RA, Dec) = (83.633083, 22.014500) (JBE position)
  • DE430 ephemeris, (RA, Dec) = (83.63311445608998, 22.01448713834) (ICRS position)
  • DE440 ephemeris, (RA, Dec) = (83.63311445608998, 22.01448713834) (ICRS position)
The procedure was run on sciserver (M. Taghizadeh-Popp et al., 2020)
A.9 NinjaSat
@@ A.10 NuSTAR
For a sample of observations, we followed the general procedure used to process Mode-06 data (e.g. D. J. Walton et al. 2016), splitting the intervals with different CHU combinations through the nusplitsc tool. From both Mode-01 and Mode-06 data, we selected data from a 70′′region around the centroid of the PSF (which, again, can move from the expected position in Mode-06 data). The selection was done through Python scripts that analyzed the images produced in each data mode interval looking for the maximum of the PSF. Again, a more precise modeling of the PSF is not needed in this case, being interested only on the event arrival times and a rough energy measurement for each photon (and not a precise spectral modeling with PSF corrections), and being the Crab by far the most luminous source in the field of view. The whole procedure was automated through the heasarc_retrieve_pipeline script v.0.2 (M. Bachetti, 2025c), installed on Sciserver (M. Taghizadeh-Popp et al., 2020). We later verified by eye that the selected regions were at least roughly consistent with the expected PSF positions.
@@
A.11 Suzaku
We downloaded the final archival products of public Crab data for Suzaku HXD PIN and GSO from HEASARC. Subsequently, we performed barycentric correction on the cleaned event FITS files using the aebarycen tool, as described in Y. Terada et al. (2008):
@@ A.12 Swift/XRT
For each observation, source events with grades 0–2 were extracted from a rectangular region (40 pixels wide) centered on the pixel with the highest count. This region encompasses approximately 94% of the Point Spread Function (PSF) of the Crab pulsar, which is located near the center of the Nebula. The event arrival times were referenced to the Solar System Barycenter (SSB) using barycorr with the Crab pulsar’s coordinates from the Jodrell Bank Monthly Ephemeris (RA = 05h34m31.972s, Dec = $22^{\circ}00^{\prime}52^{\prime\prime}.07$; A. G. Lyne et al. 1993), the JPL DE200 solar system ephemeris. The latter also accounts for time drift during the observation, using the mission clockfile (swclockcor20041120v160.fits, updated to May 6, 2024), and the spacecraft orbit as specified in the ObsID-specific attorb file.
A.13 XMM-Newton
@@
The cleaned event lists are then barycentrically corrected using the SAS task, barycen
barycen withtable=yes table=’bary.ds:EVENTS’ timecolumn=’TIME’ withsrccoordinates=yes srcra=’83.633216667’ srcdec=’22.014463889’ processgtis=yes time=0 ephemeris=DE200)
@@
A.14 XRISM
We utilized private observations of the Crab with XRISM Resolve obtained during the performance verification phase and the first cycle of the guest observation program in 2024. The processing and calibration versions of the dataset were specified as TLM2FITS= ’004_002.15Oct2023_Build7.011’, PROCVER = ’03.00.011.008’, and CALDBVER= ’gen20240315_xtd20240315_rsl20240315’, with ground-based calibration parameters incorporated for timing assignment. We then reprocessed the unscreened event file with the pre-public timing-coefficient CALDB based on in-orbit calibration and employed standard screening criteria to get cleaned event file. High- and medium-resolution grades (Hp, Mp, and Ms) were extracted by applying ”ITYPE=0:2” in xselect to the cleaned event file. A barycentric correction was then conducted on the DE200 solar system ephemeris using the command prescribed in Y. Terada et al. (2025):
@@ A.14 XRISM
Appendix B Example of re-fitted timing model
We report below the result of model_200.compare(model_430) in PINT, where model_200 is the original JBE model from the CGRO solution, and the model_430 is the result of the re-fitting of the model using the position and ephemeris used to barycenter the data, following the procedure in Section 3. The parameters marked with asterisks are those that were fitted, while the others were kept fixed to the values in the original JBE model. The apparent large differences in F0, F1, and F2 (compared to the quoted uncertainties) are due to the very low scatter of simulated data compared to the deviations produced by the different reference frames and ephemerides. TZRMJD was fixed to a TOA close to the middle of the observation span.
PARAMETER                        Crab_DE200.par   nu11002303004B_src1_de430.par   Diff_Sigma1   Diff_Sigma2
@@ Appendix B Example of re-fitted timing model
START                                   60341.0           60368.999999907888043
FINISH                                  60370.0            60402.00000009055646
TRES                                      27.06             0.09687254051175789                             !
RAJ                              5h34m31.97232s                  5h34m31.97232s
DECJ                              22d00m52.069s                 22d00m52.06896s
@@
F0                          29.5638899201697(0)             29.5638898975506(4)                   -56643.62 ! *
PEPOCH                                  60356.0                         60356.0
F1                 -3.6672999999999996(0)10-10           -3.667070990(4)10-10                    51907.78 ! *
F2                               -1.99(0)10-20                -3.475(18)10-22                    10868.98 ! *
TZRMJD                          60356.000000369           60386.027999965299387                             !
@@
SEPARATION                      0.000040 arcsec

### v3

@@ B.8 NICER
  • DE200 ephemeris, (RA, Dec) = (83.633083, 22.014500) (FK5)
  • DE430 ephemeris, (RA, Dec) = (83.63311445608998, 22.01448713834) (ICRS)
  • DE440 ephemeris, (RA, Dec) = (83.63311445608998, 22.01448713834) (ICRS)
The procedure was run on sciserver (Taghizadeh-Popp et al., 2020)
B.9 NinjaSat
@@ B.10 NuSTAR
For a sample of observations, we followed the general procedure used to process Mode-06 data (e.g. Walton et al. 2016), splitting the intervals with different CHU combinations through the nusplitsc tool. From both Mode-01 and Mode-06 data, we selected data from a 70′′ region around the centroid of the PSF (which, again, can move from the expected position in Mode-06 data). The selection was done through Python scripts that analyzed the images produced in each data mode interval looking for the maximum of the PSF. Again, a more precise modeling of the PSF is not needed in this case, being interested only on the event arrival times and a rough energy measurement for each photon (and not a precise spectral modeling with PSF corrections), and being the Crab by far the most luminous source in the field of view. The whole procedure was automated through the heasarc_retrieve_pipeline script v.0.2 (Bachetti, 2025a), installed on Sciserver (Taghizadeh-Popp et al., 2020). We later verified by eye that the selected regions were at least roughly consistent with the expected PSF positions.
@@
B.11 Suzaku
We downloaded the final archival products of public Crab data for Suzaku HXD PIN and GSO from HEASARC. Subsequently, we performed barycentric correction on the cleaned event FITS files using the aebarycen tool, as described in Terada et al. (2008a):
@@ B.12 Swift/XRT
For each observation, source events with grades 0–2 were extracted from a rectangular region (40 pixels wide) centered on the pixel with the highest count. This region encompasses approximately 94% of the Point Spread Function (PSF) of the Crab pulsar, which is located near the center of the Nebula. The event arrival times were referenced to the Solar System Barycenter (SSB) using barycorr with the Crab pulsar’s coordinates from the Jodrell Bank Monthly Ephemeris (RA = 05h34m31.972s, Dec = $22^{\circ}00^{\prime}52^{\prime\prime}.07$; Lyne et al. 1993), the JPL-DE200 solar system ephemeris. The latter also accounts for time drift during the observation, using the mission clockfile (swclockcor20041120v160.fits, updated to May 6, 2024), and the spacecraft orbit as specified in the ObsID-specific attorb file.
B.13 XMM-Newton
@@
The cleaned event lists are then barycentrically corrected using the SAS task, barycen, and the FK5 (R.A., Dec.) 05:34:31.972 +22:00:52.070:
barycen withtable=yes table=’bary.ds:EVENTS’ timecolumn=’TIME’ withsrccoordinates=yes
srcra=’83.633216667’ srcdec=’22.014463889’ processgtis=yes time=0 ephemeris=DE200)
@@
B.14 XRISM
We utilized private observations of the Crab with XRISM Resolve obtained during the performance verification phase and the first cycle of the guest observation program in 2024. The processing and calibration versions of the dataset were specified as TLM2FITS= ’004_002.15Oct2023_Build7.011’, PROCVER = ’03.00.011.008’, and CALDBVER= ’gen20240315_xtd20240315_rsl20240315’, with ground-based calibration parameters incorporated for timing assignment. We then reprocessed the unscreened event file with the pre-public timing-coefficient CALDB based on in-orbit calibration and employed standard screening criteria to get cleaned event file. High- and medium-resolution grades (Hp, Mp, and Ms) were extracted by applying "ITYPE=0:2" in xselect to the cleaned event file. A barycentric correction was then conducted on the DE200 solar system ephemeris using the command prescribed in Terada et al. (2025):
@@ B.14 XRISM
B.15 XTE
We obtained a list of all RXTE/PCA observations pointed less than 0.1 degrees away from the Crab pulsar by querying the xtemaster table at HEASARC, using the astroquery.heasarc module. We selected only observations containing data in science event modes. For each of these observations, we downloaded locally the event data (fname) and the relevant spacecraft file (orbit_file), and barycentered the event data using the command
barycorr fname outfile ra=83.63311446 dec=22.01448714 ephem=JPLEPH.DE440 refframe=ICRS
orbitfiles=orbit_file clobber=yes
To save space, we did not save the full information of the barycentered event files, but only the arrival times of the photons and the PHA column that could be used to extract energy-selected event lists. We did not perform an energy selection for this work.
We processed in total 330 observations between MJD 50205–55927 (1996–2012). The procedure was executed on the Fornax infrastructure at NASA [^22] https://pcos.gsfc.nasa.gov/Fornax/ (Jaffe, 2024).
Appendix C Example of re-fitted timing model
We report below the result of model_200.compare(model_405) in PINT, where model_200 is the original JBE model from the CGRO solution, closest to the start of the observation (MJD 55427.9964794566), and the model_405 is the result of the re-fitting of the model using the position and ephemeris used to barycenter the data, following the procedure in Section III. The parameters marked with asterisks are those that were fitted, while the others were kept fixed to the values in the original JBE model. The apparent large differences in F0, F1, and F2 (compared to the quoted uncertainties) are due to the very low scatter of simulated data (see the model TRES) compared to the deviations produced by the different reference frames and ephemerides. TZRMJD was fixed to a TOA close to PEPOCH.
PARAMETER          Crab_55427.9964794566_DE200.par     crab_12228_none.par   Diff_Sigma1   Diff_Sigma2
@@ Appendix C Example of re-fitted timing model
START                                      55409.0   55408.000000064643586
FINISH                                     55440.0   55441.000000165256875
DILATEFREQ                                   False                   False
DMDATA                                       False                   False
NTOA                                             0                     501
TRES                                         40.38     0.09930508316423779                             !
POSEPOCH                                   40706.0                 40706.0
PX                                             0.0                     0.0
RAJ                                 5h34m31.97232s          5h34m31.99992s
DECJ                                 22d00m52.069s          22d00m52.0992s
@@
F0                              29.721114704239(0)     29.7211147070001(4)                     6878.81 ! *
PEPOCH                                     55424.0                 55424.0
F1                   -3.7117799999999996(0)10E-18  -3.711789115(4)10E-18                    -2055.94 ! *
F2                    2.4400000000000004(0)10E-20      2.42957(18)10E-20                      -57.71 ! *
PLANET_SHAPIRO                               False                   False
DM                                             0.0                     0.0
TZRMJD                             55424.000000163   55424.037999928536006                             !
@@
PHOFF                                       0.0(0)          0.0(2.0)10E-7             nan          0.00 *
CHI2                                       Missing      494.06112706703493
CHI2R                                      Missing      0.9960909819899897
SEPARATION                         0.385001 arcsec